# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v4

@@ B.2 Probit Model
Notes: Probit model with $\varepsilon_{it}\sim\mathcal{N}(0,1)$. Design 2: $\{\alpha_{k,i}^{0},\gamma_{k,t}^{0}\}\sim\mathcal{N}(0,1/4)$, $X_{it}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0}+\gamma_{1,t}^{0}+\gamma_{2,t}^{0},1)$, and $X_{i0}\sim\mathcal{N}(\alpha_{1,i}^{0}+\alpha_{2,i}^{0},1)$. The null hypothesis $H_{0}:(\beta_{0},\rho_{0})^{\prime}=(0.5+\delta,0.5+\delta)^{\prime}$ for the dynamic model, and $H_{0}:\beta_{0}=0.5+\delta$ for the static model, with a significance level $5\%,$ where $\delta\in\{\pm 0.2,\pm 0.15,\pm 0.1\}$. The number of replications is $1000$.
C Additional Empirical Results
C.1 Sample
@@
The data set is constructed from waves $20-30$ of PSID, covering the survey years 1987 to 1997. We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999. Each survey wave includes questions about employment status in the prior calendar year, such as, ‘About how many weeks did (he/she) work on that job last year?’ and ‘During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?’ We use responses to these questions to construct individual labor force participation (LFP) status for the prior year. Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on. In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
We also restrict our sample to individuals whose LFP status changed at least once during the sample period. Individuals whose labor supply status never changed during the sample period are not informative for our analysis, because the dependent variable could be perfectly predicted by setting the individual fixed effects in the intercept of our models to plus or negative infinity.
The final sample consists of a panel of $N=90$ individuals over $T=10$ years. Table C.1 presents the summary statistics. Figure C.1 shows the average percentage of individuals with given current and past LFP status. Figure C.2 plots the LFP transition rates by year.
Table C.1: Descriptive Statistics
@@
Labor Force Participation 0.482 0.500 0.000 1.000
Lagged Participation 0.458 0.498 0.000 1.000
Number of Children 2.24 1.098 1.000 6.000
  • Notes: Summary statistics are computed based on a sample of 90 individuals observed from 1987 to 1996. All variables have 900 observations. Data source: PSID 1987 - 1997.
Figure C.1: Labor Force Participation Contingency Table. Each cell contains the percentage of individuals with given current and past labor force participation status.
Figure C.2: Labor Force Participation Status Transition Rates By Year
C.2 Extra Empirical Results
Table C.2: Estimated Heterogeneous Effects on Labor Force Participation

### v5

@@ F.3 Probit and Logit under Design 2
Table F.12: Empirical Size and Power from LR Test for Logit and Probit Model, Design 2
Size Power at $\delta$
$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $47$ 2020 $79$ $45$ $12$ $26$
$(60,60)$ $43$ $10$ $100$ $73$ $40$ $93$
$(90,90)$ $42$ 99 $100$ $94$ $81$ $100$
$(120,120)$ $40$ 77 $100$ $99$ $97$ $99$
$(150,150)$ $39$ 55 $100$ $100$ $100$ $100$
Static Probit Model
$(30,30)$ $77$ $39$ $97$ $77$ $14$ $15$
$(60,60)$ $76$ 2020 $100$ $96$ $34$ $83$
$(90,90)$ $72$ $14$ $100$ $99$ $85$ $96$
$(120,120)$ $72$ $10$ $99$ $99$ $98$ $98$
$(150,150)$ $72$ 99 $99$ $99$ $99$ $99$
Dynamic Logit Model
$(30,30)$ $43$ $18$ $75$ $41$ 1919 $35$
$(60,60)$ $38$ 99 $100$ $69$ $44$ $96$
$(90,90)$ $34$ 77 $100$ $93$ $86$ $100$
$(120,120)$ $33$ 66 $100$ $99$ $98$ $100$
$(150,150)$ $31$ 55 $100$ $100$ $100$ $100$
Dynamic Probit Model
$(30,30)$ $71$ $37$ $95$ $74$ $25$ $38$
$(60,60)$ $66$ 1919 $100$ $93$ $47$ $96$
$(90,90)$ $63$ $14$ $100$ $100$ $90$ $100$
$(120,120)$ $64$ $11$ $99$ $99$ $99$ $99$
$(150,150)$ $62$ 88 $99$ $99$ $99$ $99$
[…]
Table F.13: Empirical Size and Power from LM Test for Logit and Probit Model, Design 2
Size Power at $\delta$
$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $42$ $18$ $77$ $42$ $12$ $28$
$(60,60)$ $42$ 99 $99$ $72$ $41$ $93$
$(90,90)$ $41$ 99 $100$ $94$ $82$ $100$
$(120,120)$ $40$ 77 $100$ $99$ $97$ $100$
$(150,150)$ $39$ 55 $100$ $99$ $100$ $100$
Static Probit Model
$(30,30)$ $69$ $34$ $96$ $73$ $12$ $17$
$(60,60)$ $74$ $18$ $100$ $96$ $37$ $76$
$(90,90)$ $71$ $14$ $100$ $99$ $86$ $88$
$(120,120)$ $72$ $10$ $99$ $99$ $96$ $91$
$(150,150)$ $72$ 99 $99$ $100$ $98$ $94$
Dynamic Logit Model
$(30,30)$ $38$ $16$ $73$ $38$ 1919 $36$
$(60,60)$ $37$ 99 $100$ $68$ $45$ $94$
$(90,90)$ $34$ 77 $100$ $93$ $86$ $99$
$(120,120)$ $33$ 66 $100$ $99$ $98$ $99$
$(150,150)$ $31$ 55 $100$ $100$ $100$ $99$
Dynamic Probit Model
$(30,30)$ $59$ $30$ $94$ $70$ 2121 $38$
$(60,60)$ $62$ $17$ $100$ $92$ $48$ $91$
$(90,90)$ $60$ $13$ $100$ $99$ $90$ $94$
$(120,120)$ $62$ $10$ $99$ $99$ $97$ $96$
$(150,150)$ $62$ 77 $99$ $99$ $99$ $96$
[…]
Table F.14: Empirical Size and Power from Wald Test for Logit and Probit Model, Design 2
Size Power at $\delta$
$(N,T)$ Uncorrected Analytical 0.2-0.2 0.1-0.1 $0.1$ $0.2$
Static Logit Model
$(30,30)$ $32$ $15$ $75$ $39$ $10$ $29$
$(60,60)$ $41$ 99 $100$ $72$ $41$ $95$
$(90,90)$ $42$ 99 $100$ $94$ $81$ $100$
$(120,120)$ $40$ 77 $100$ $99$ $97$ $100$
$(150,150)$ $39$ 55 $100$ $100$ $100$ $100$
Static Probit Model
$(30,30)$ $37$ $29$ $91$ $68$ $10$ 1919
$(60,60)$ $71$ $18$ $99$ $95$ $39$ $92$
$(90,90)$ $72$ $14$ $99$ $99$ $88$ $99$
$(120,120)$ $72$ $10$ $99$ $99$ $98$ $99$
$(150,150)$ $72$ 99 $99$ $99$ $99$ $99$
Dynamic Logit Model
$(30,30)$ $27$ $15$ $69$ $34$ $18$ $36$
$(60,60)$ $36$ 88 $100$ $68$ $45$ $97$
$(90,90)$ $34$ 77 $100$ $92$ $86$ $100$
$(120,120)$ $33$ 66 $100$ $99$ $98$ $100$
$(150,150)$ $31$ 55 $100$ $100$ $100$ $100$
Dynamic Probit Model
$(30,30)$ $25$ $26$ $88$ $60$ 2121 $39$
$(60,60)$ $56$ $16$ $100$ $91$ $49$ $98$
$(90,90)$ $60$ $13$ $99$ $99$ $91$ $100$
$(120,120)$ $64$ $10$ $99$ $99$ $98$ $98$
$(150,150)$ $62$ 77 $99$ $99$ $99$ $99$
[…]
Table F.15: Comparisons of Uncorrected, Corrected, and Jackknife Estimates for Logit and Probit, Design 2
$(N,T)$ $(30,30)$ $(60,60)$ $(90,90)$
Bias RMSE Bias RMSE Bias RMSE
Static Logit Model
$\widehat{\beta}$ $0.400$ $1.435$ $0.101$ $0.125$ $0.060$ $0.072$
$\widehat{\beta}_{L}$ $0.093$ $0.188$ $0.022$ $0.056$ $0.010$ $0.036$
$\widehat{\beta}_{J}$ 2.314-2.314 $4.328$ 0.172-0.172 $0.406$ 0.031-0.031 $0.070$
Static Probit Model
$\widehat{\beta}$ $0.846$ $2.168$ $0.151$ $0.261$ $0.074$ $0.082$
$\widehat{\beta}_{L}$ $0.182$ $0.339$ $0.041$ $0.068$ $0.019$ $0.033$
$\widehat{\beta}_{J}$ 4.929-4.929 $14.383$ 0.504-0.504 $0.981$ 0.105-0.105 $0.214$
Dynamic Logit Model
$\widehat{\beta}$ $0.415$ $0.858$ $0.104$ $0.137$ $0.061$ $0.074$
$\widehat{\beta}_{L}$ $0.104$ $0.201$ $0.023$ $0.060$ $0.012$ $0.043$
$\widehat{\beta}_{J}$ 2.442-2.442 $4.223$ 0.231-0.231 $0.481$ 0.043-0.043 $0.105$
$\widehat{\rho}$ 0.120-0.120 $0.257$ 0.055-0.055 $0.108$ 0.036-0.036 $0.071$
$\widehat{\rho}_{L}$ 0.047-0.047 $0.205$ 0.016-0.016 $0.089$ 0.010-0.010 $0.059$
$\widehat{\rho}_{J}$ $0.020$ $0.218$ $0.003$ $0.092$ $0.002$ $0.060$
Dynamic Probit Model
$\widehat{\beta}$ $1.040$ $2.343$ $0.184$ $0.287$ $0.080$ $0.090$
$\widehat{\beta}_{L}$ $0.209$ $0.332$ $0.052$ $0.082$ $0.020$ $0.049$
$\widehat{\beta}_{J}$ 3.352-3.352 $7.708$ 0.558-0.558 $1.125$ 0.118-0.118 $0.217$
$\widehat{\rho}$ 0.046-0.046 $0.169$ 0.022-0.022 $0.072$ 0.014-0.014 $0.046$
$\widehat{\rho}_{L}$ 0.028-0.028 $0.145$ 0.011-0.011 $0.067$ 0.009-0.009 $0.053$
$\widehat{\rho}_{J}$ 0.029-0.029 $0.169$ 0.006-0.006 $0.068$ 0.002-0.002 $0.044$
[…]
Notes: Simulated separately from Table F.11. For dynamic models, We set $\tau=1$ in $L\left(\theta\right)$, producing corrected estimators $\widehat{\rho}_{L}$ and $\widehat{\beta}_{L}$ for $\rho_{0}$ and $\beta_{0}$ respectively.
F.4 Two-way Fixed-Effect Probit
In this section, we consider a dynamic probit model with two-way scalar fixed effects appearing only in the intercept. This is a special case of the two-way DN-HP model. We take Design 1 in Fernández-Val and Weidner (2016) and simulate our method, the jackknife, and the bootstrap here. We set $N=56$ and $T\in\left\{14,28,56\right\}$, and consider
$Y_{it}=\mathds{1}\left\{\rho_{0}Y_{it-1}+\beta_{0}Z_{it}+\alpha_{i}^{0}+\gamma_{t}^{0}+\varepsilon_{it}>0\right\}$
with $Y_{i0}=\mathds{1}\left\{\beta_{0}Z_{i0}+\alpha_{i}^{0}+\gamma_{0}^{0}+\varepsilon_{i0}>0\right\}$, where $\rho_{0}=0.5$, $\beta_{0}=1$, $\varepsilon_{it}$ and $\varepsilon_{i0}$ are standard normal, $\left\{\alpha_{i}^{0},\gamma_{t}^{0},\gamma_{0}^{0}\right\}\sim\mathcal{N}\left(0,0.0625\right)$, and $Z_{it}=Z_{it-1}/2+\alpha_{i0}+\gamma_{t0}+\nu_{it}$ with $\nu_{it}\sim\mathcal{N}\left(0,1/2\right)$ and $Z_{i0}\sim\mathcal{N}\left(0,1\right)$. We set $\tau\in\left\{1,2\right\}$. For the hypothesis tests, we only present the results from the LR tests, using the uncorrected likelihood $\widehat{l}\left(\theta\right)$, the corrected likelihood $L\left(\theta\right)$ (setting $\tau=1$), and the bootstrap. We set the number of repetitions to $499$ for the bootstrap and run $1000$ replications. The rest are the same as in Section 4.
Table F.16 presents the simulation results. First, the biases and RMSEs of $\widehat{\beta}$, $\widehat{\beta}_{J}$, $\widehat{\rho}$, and $\widehat{\rho}_{J}$ are in line with the original simulations of Fernández-Val and Weidner (2016). For instance for $\left(N,T\right)=\left(56,14\right)$, the bias of $\widehat{\beta}$ is $0.182$, roughly $18\%$ relative to the true value of $1$, and its RMSE is $0.221$, roughly $22\%$. For $\widehat{\beta}_{J}$, the bias is $8\%$ and the RMSE is $11\%$. These are very close to the results of Fernández-Val and Weidner (2016). As for the corrected estimators, we find that the bootstrap estimators $\widehat{\beta}_{B}$ and $\widehat{\rho}_{B}$ are generally the best, outperforming the jackknife estimators $\widehat{\beta}_{J}$ and $\widehat{\rho}_{J}$ and our corrected estimators $\widehat{\beta}_{L}^{\left(\tau\right)}$ and $\widehat{\rho}_{L}^{\left(\tau\right)}$. The biases of $\widehat{\beta}_{B}$ and $\widehat{\rho}_{B}$ are also comparable to the parameter-based procedure of Fernández-Val and Weidner (2016). Our procedure outperforms the jackknife in estimating $\beta_{0}$ (but not $\rho_{0}$) when the sample size is small. However, the differences between all estimators (except the MLE) become small as the sample size increases. While all being analytical corrections, our corrected estimator performs slightly worse than the parameter-based corrected estimator of Fernández-Val and Weidner (2016). This can be anticipated, because our procedure produces the bias correction effect only indirectly via the likelihood, whereas the parameter-based approach applies directly to the estimator. However, we would like to point out that the procedure of Fernández-Val and Weidner (2016) does not apply to the DN-HP model without modifications.
Table F.16: Comparisons of Uncorrected and Corrected Two-Way Fixed-Effect Probit Estimates
$(N,T)$ $(56,14)$ $(56,28)$ $(56,56)$
Bias RMSE Bias RMSE Bias RMSE
$\widehat{\beta}$ $0.182$ $0.221$ $0.093$ $0.120$ $0.062$ $0.078$
$\widehat{\beta}_{L}^{(1)}$ $0.063$ $0.156$ $0.020$ $0.073$ $0.012$ $0.047$
$\widehat{\beta}_{L}^{(2)}$ $0.076$ $0.159$ $0.022$ $0.075$ $0.012$ $0.047$
$\widehat{\beta}_{J}$ 0.083-0.083 $0.151$ 0.028-0.028 $0.081$ 0.007-0.007 $0.047$
$\widehat{\beta}_{B}$ 0.045-0.045 $0.105$ 0.017-0.017 $0.069$ 0.002-0.002 $0.045$
$\widehat{\rho}$ 0.226-0.226 $0.266$ 0.098-0.098 $0.134$ 0.047-0.047 $0.081$
$\widehat{\rho}_{L}^{(1)}$ 0.100-0.100 $0.166$ 0.030-0.030 $0.090$ 0.016-0.016 $0.065$
$\widehat{\rho}_{L}^{(2)}$ 0.102-0.102 $0.171$ 0.024-0.024 $0.091$ 0.011-0.011 $0.065$
$\widehat{\rho}_{J}$ $0.046$ $0.161$ $0.015$ $0.095$ $0.000$ $0.067$
$\widehat{\rho}_{B}$ $0.015$ $0.131$ $0.010$ $0.088$ 0.001-0.001 $0.064$
[…]
Table F.17 presents the empirical sizes and powers of the LR test using the uncorrected likelihood (size only), our corrected likelihood, and the bootstrap. Briefly speaking, our procedure delivers an empirical size close to the nominal level of $5\%$ and provide a reasonable empirical power, starting from $T=28$. On the opposite, the empirical size from the bootstrap is close to $5\%$ for all $T\in\left\{14,28,56\right\}$, but its power is relatively low, especially near the true hypothesis.
Table F.17: Empirical Size and Power from LR Test for Two-Way Fixed-Effect Probit Estimates
(a) Empirical Size
Size
$(N,T)$ Uncorrected Analytical Bootstrap
$(56,14)$ $53$ $17$ 44
$(56,28)$ $30$ 88 33
$(56,56)$ $25$ 77 55
(b) Empirical Power
Power at $\delta$ (Analytical) Power at $\delta$ (Bootstrap)
$(N,T)$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$ 0.2-0.2 0.1-0.1 $0.1$ $0.2$
$(56,14)$ $79$ $37$ $33$ $70$ $35$ $10$ 66 1919
$(56,28)$ $97$ $46$ $45$ $95$ $87$ $25$ $10$ $56$
$(56,56)$ $100$ $79$ $72$ $100$ $100$ $70$ $27$ $94$
[…]
G Additional Empirical Results
G.1 Sample
@@
The data set is constructed from waves 20–30 of PSID, covering the survey years 1987 to 1997. We selected this time span because 1997 was the last year PSID conducted annual surveys, switching to a biennial schedule in 1999. Each survey wave includes questions about employment status in the prior calendar year, such as, “About how many weeks did (he/she) work on that job last year?” and “During the weeks that (he/she) worked, about how many hours did (he/she) usually work per week?” We use responses to these questions to construct individual labor force participation (LFP) status for the prior year. Thus, wave 30 (1997) provides LFP data for 1996, wave 29 (1996) for 1995, and so on. In our final sample, labor force participation status is observed from 1987 to 1996, with lagged participation status observed from 1986 to 1995.
We restrict the data to the informative sample of individuals whose current and lagged LFP status each varied at least once during the sample period. This ensures within-person variation in both variables, which is necessary for identification of the dynamic model. The final sample contains $N=86$ single mothers observed for $T=10$ consecutive years. Table G.1 reports the summary statistics. Figure G.1 displays the joint distribution of current and lagged LFP status, and Figure G.2 plots the year-specific transition rates between employment and non-employment.
Table G.1: Descriptive Statistics
@@
Labor Force Participation 0.491 0.500 0.000 1.000
Lagged Participation 0.467 0.499 0.000 1.000
Number of Children 2.25 1.104 1.000 6.000
  • Notes: Summary statistics are computed based on a sample of 86 individuals observed from 1987 to 1996. All variables have 860 observations. Data source: PSID 1987–1997.
Figure G.1: Labor Force Participation Contingency Table
[…]
Notes: Each cell contains the percentage of individuals with given current and past labor force participation status.
Figure G.2: Labor Force Participation Status Transition Rates By Year
G.2 Extra Empirical Results
Table G.2: Estimated Heterogeneous Effects on Labor Force Participation